Given the following hypotheses: H0: μ = 540 H1: μ ≠ 540 A random sample of 10 observations is selected from a normal population. The sample mean was 550 and the sample standard deviation was 6. Using the 0.10 significance level: State the decision rule. (Negative amount should be indicated by a minus sign. Round your answers to 3 decimal places.) Compute the value of the test statistic. (Round your answer to 3 decimal places.)
Answer #1

ANSWER:

Given that,

Sample size = n = 10

Sample mean = = 550

Standard deviation = s = 6

a) Null and alternative hypothesis is

H0: μ = 540

H1: μ ≠ 540

Test statistic is

=550-540/(6/3.1623)

= 10/1.897

=5.271

Degrees of freedom = n - 1 = 10 - 1 = 9

Critical value = 1.833 ( Using t table)

Decision Rule : Test statistic > critical value we reject null hypothesis.